OpenStudy (anonymous):

rewrite log4+log2-log5 as a single logarithm

mathslover (mathslover):

Use the following properties of logarithm : \(\log a + \log b = \log (ab) \) and \(\log a - \log b = \log \left(\cfrac{a}{b}\right) \)